OpenStudy (anonymous):

Verify the identity. tan (x + pi/2) = -cot x please show your work! :)

OpenStudy (anonymous):

Do you know the sum rule for tangent?

OpenStudy (anonymous):

i have it in front of me but im not sure what it means

OpenStudy (anonymous):

ok... the sum rule is: \(\large tan(A+B)=\frac{tanA+tanB}{1-tanA \cdot tanB} \) so let A = x, B = \(\frac{\pi}{2} \) and plug it in to the formula and simplify..

OpenStudy (anonymous):

so would it be: tan (x+π/2) = (tan x + tan π/2)/ (1-tan x)(tan π/2) @ByteMe

OpenStudy (anonymous):

yes... but i'm thinking there's a problem here because tan(pi/2) is undefined... It is better to rewrite it like this: \(\large tan(A+B)=\frac{sin(A+B)}{cos(A+B)} \) = that one ^^^ will work....

OpenStudy (anonymous):

so that would b my final answer? @ByteMe

OpenStudy (anonymous):

no... you're final answer is working to -cotx... you need to simplify it so that you end up with -cotx.

OpenStudy (anonymous):

tan(x+π/2) = sin(x+π/2)/cos(x+π/2)

OpenStudy (anonymous):

could u show me how?

OpenStudy (anonymous):

yes... ^^^ now simplify that using sum formulas for sine and cosine..

OpenStudy (anonymous):

sin(A+B) = sinAcosB + sinBcosA cos(A+B) = cosAcosB - sinAsinB again let A=x, and B=pi/2 and then simplify...

OpenStudy (anonymous):

then how do i get to -cot x from there

OpenStudy (anonymous):

sin(A+B) = sin(x + pi/2) = sinx cos(pi/2) + sin(pi/2) cos x = cos x

OpenStudy (anonymous):

can you do the cosine part?

OpenStudy (anonymous):

yeah one sec

OpenStudy (anonymous):

cos (A+B) = cos(x+pi/2) = cos x cos(pi/2) - sin x sin(pi/2)

OpenStudy (anonymous):

@ByteMe

OpenStudy (anonymous):

that's correct... :)

OpenStudy (anonymous):

notice that cos(pi/2) = 0, sin(pi/2) = 1 substitute those and simplify... wat u got????

OpenStudy (anonymous):

sin x +cos x and cos x - sin x

OpenStudy (anonymous):

cos(x + pi/2) = cosx cos(pi/2) - sinx sin(pi/2) = cosx (0) - sinx (1) = 0 - sinx = -sinx and since sin(x + pi/2) = cosx, then tan(x + pi/2) = sin(x + pi/2) / cos(x + pi/2) = cosx / -sinx = - cosx/sinx = -cotx

OpenStudy (anonymous):

k???

OpenStudy (anonymous):

that makes sense. could you help me with one more plz?!?!

OpenStudy (anonymous):

sure... just post it up as a separate question... :)

OpenStudy (anonymous):

close this question and post a new question...

OpenStudy (anonymous):

k